OpenStudy (pulsified333):

How much 90 percent vinegar do you have to add to a gallon of 5 percent vinegar to get 20 percent vinegar? Hint: You want the final percentage of pure vinegar to be 20%. You have 1+x gallons of fluid where x is the number of gallons you add. Think about how much pure vinegar is in those 1+x gallons.

OpenStudy (pulsified333):

@Directrix

Directrix (directrix):

From: @stankytukus 0.90x + 0.05*1 = 0.20(x+1) where x is the amount of 90% vinegar Multply through by 100 90x + 5 = 20x + 20 70x = 15 @stankytukus requests that you solve for x @Pulsified333

OpenStudy (pulsified333):

x=15/70?

Directrix (directrix):

Yes. And, as a decimal, how many gallons is that?

OpenStudy (pulsified333):

0.214285714

Directrix (directrix):

Check to see if your answer has to be given to a nearest decimal place. The exact answer is 15/70 which equals 3/14 after dividing out the common factor.

OpenStudy (pulsified333):

it is correct :D

Directrix (directrix):

Good to hear.